1. The area of a trapezoid is 1/2 ah + 1/2 bh. Express the area as product of two factors.
1
Expert's answer
2014-07-23T11:09:37-0400
Expression of 1/2 ah + 1/2 bh=1/2*(ah+bh)=h/2*(a+b) as product of factors.